# MLLM4PUE: Toward Universal Embeddings in Computational Pathology through Multimodal LLMs

### v1

@@ 5 Results
Table 3: Zero-shot composed retrieval performance on the Quilt-VQA and Quilt-VQA-RED datasets, with Recall metrics reported at various thresholds. MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [16], PathCap [38], Quilt1M [17], respectively. Bold values indecate the best performance.
Prompt-tuning Prompt Arch-PubMed [13] Arch-book [13]
i2t t2i i2t t2i
R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10
no no prompt 0.086 0.135 0.087 0.137 0.147 0.213 0.149 0.226
no “in one word” 0.092 0.137 0.095 0.142 0.152 0.236 0.158 0.233
yes no prompt 0.091 0.139 0.091 0.141 0.152 0.233 0.161 0.246
no “in one word” 0.094 0.143 0.102 0.149 0.157 0.243 0.165 0.258
MLLM4PUE 0.098 0.155 0.109 0.162 0.169 0.253 0.177 0.274
Table 4: Impact of prompt-tuning on two retrieval datasets. All models are trained on the Openpath dataset [16] for 1000 steps. Task i2t and t2i denote image-to-text and text-to-image retrieval. R@5 and R@10 represent Recall@5 and Recall@10, respectively. Bold values indicate the best performance.
@@
5.1 Zero-shot Cross-modal Retrieval
We first evaluate our methods on the retrieval task, where the query is an image or text, with the target being the corresponding paired text or image. Recall@k metric is employed to assess the model’s ability to retrieve relevant pathology image-text pairs among top k results. Our MLLM4PUE models are pretrained on Openpath [16], PatchCap [38], and Quilt1M [17]. As shown in Table 2, our models consistently outperform the baseline models, especially at higher recall values. Notably, in the image-to-text (i2t) task on Arch-book dataset [13], our MLLM4PUE trained on Openpath dataset [16] achieves a Recall@5 of 0.185, indicating an approximate 9% improvement over PLIP [16]. Moreover, our PathCap-pretrained model surpasses the performance of PathCLIP [38] on both benchmark datasets. A similar enhancement is observed in the text-to-image (t2i) retrieval task. For instance, on the Arch-PubMed dataset [13], our model achieves a Recall@10 of 0.193, marking a 10% increase in performance. These results collectively demonstrate that our framework excels at capturing multimodal embeddings more effectively than CLIP-based models with the same training data. Furthermore, it is important to note that the E5-V model [19], although sharing the same underlying architecture as ours, yields significantly lower recall scores across all tasks and datasets due to the lack of fine-tuning with pathology-specific data.
5.2 Zero-shot Patch Classification
We then evaluate the performance of zero-shot classification on patch-level pathology images. Specifically, each query is an image, and the target is an extended sentence of the class name, constructed using a template outlined in sec 3.4. Following CONCH [31], we conduct 100 experimental trials for each dataset, randomly selecting one template from a set of 21 templates for each trial. The detailed descriptions are available in the Supplementary Material. We present the distribution of weighted F1 scores (wF1) from these 100 experiments in Fig. 4, where the top, middle, and bottom lines of each box depict the first quartile, median, and third quartile of the wF1 score distribution, respectively. Each point in the figure indicates a single template’s performance.
Fig. 4 illustrates that our MLLM4PUE consistently surpasses baseline models trained on identical datasets. For instance, in the “DigestpathSignet” task with the Openpath dataset [16], which focuses on identifying signet ring cells, MLLM4PUE-O achieves a median wF1 score of 0.665, marking an improvement of over 30% compared to PLIP’s median of 0.329. When the training data changes to PathCLIP [38], MLLM4PUE-P also shows considerable advancements, achieving a median score of 0.854 compared to PathCLIP’s 0.361. This implies a performance increase of roughly 50%. Similarly, using the Quiltnet dataset [17], MLLM4PUE-Q excels in the “Sicap” dataset, reaching a median wF1 of 0.402, far exceeding Quiltnet’s 0.164. In the “Osteo” dataset, MLLM4PUE-Q achieves a median wF1 score of 0.671, significantly outclassing Quiltnet with an improvement exceeding 40%. Moreover, when compared with E5V, MLLM4PUE demonstrates superior performance, delivering higher wF1 scores even with the same architectural foundation. These results confirm the superiority of MLLM4PUE in enhancing zero-shot classification performance through effective multimodal data integration, establishing it as a robust choice for various pathology image classification tasks.
5.3 Zero-shot Composed Retrieval
We further evaluate the performance of composed retrieval on two datasets, where the task involves using an image and a question as the query to retrieve the correct answer. The Recall@k metric is employed to assess how well the model can identify the correct answer within the top k candidates. As shown in Table 3, our proposed MLLM4PUE demonstrates significant improvements compared to baseline methods, even when trained with the same pre-training data. Specifically, on the Quilt-VQA dataset [35], MLLM4PUE achieves an increase in Recall@5 by approximately 4%, 3%, and 30% over PLIP [16], PathCLIP [38], and QuiltNet [17], respectively. Similarly, on the Quilt-RED-VQA dataset [35], MLLM4PUE pre-trained on Quilt1M [17], has achieved a Recall@5 of 0.675, which represents the best performance among all baseline methods. Interestingly, the E5-V model [19] exhibits superior performance compared to all CLIP-based models, despite lacking fine-tuning on pathology-specific data. This demonstrates the robust capacity of MLLM backbone in multimodal embedding fusion, highlighting its proficiency in handling diverse data inputs effectively.
5.4 Ablation studies
5.4.1 Impact of Prompts
To assess the impact of different prompt strategies, we conduct a comparative analysis between our MLLM4PUE and two alternative approaches: no prompt - where images or text are directly input into the model without prompts, and no “in one word” - employing a similar prompt to our method but omitting the phrase “in one word”. Table 4 presents the performance of these methods, both with and without fine-tuning. All models are trained on Openpath dataset [16] for 1000 steps. For models employing fine-tuning, each method is adjusted with its respective prompts. In contrast, the no fine-tuning scenario involves simply altering the prompt for the already finely-tuned MLLM4PUE model. Our method shows significant improvements on zero-shot retrieval tasks compared to the no prompt and no “in one word”. For the Arch-PubMed dataset, MLLM4PUE attains Recall@5 scores of 0.098 and 0.109 for image-to-text and text-to-image tasks, respectively, marking the best performance in those categories. Similarly, in the Arch-book dataset, MLLM4PUE achieves the top Recall@5 and Recall@10 scores, indicating superior retrieval effectiveness. These results highlight the significant enhancements provided by our prompt, emphasizing its role in facilitating more accurate zero-shot retrieval in complex datasets by better aligning multimodal inputs.
5.4.2 Fusion Method in Composed Retrieval
@@
This experiment is designed to highlight the effectiveness of our MLLM4PUE in modality fusion while exploring the limitations of CLIP in handling composed retrieval tasks. As shown in Table 5, the results reveal that QuiltNet [17], a CLIP-based model, shows little difference in performance between using only the question text and integrating image embeddings. This indicates that CLIP’s method lacks robust modality integration, as the inclusion of visual information fails to significantly enhance its effectiveness. Conversely, our model, MLLM4PUE, demonstrates substantial improvements by effectively fusing image and text information. MLLM4PUE outperforms QuiltNet in every setting. The true strength of our approach becomes evident when employing a prompt-based method, leading to significantly higher recall scores. Specifically, it achieves Recall@5 values of 0.475 on Quilt-VQA and 0.675 on Quilt-VQA-RED, representing approximately an 8% and 12% improvement over the simple addition method, respectively, and an 18% and 25% improvement over using the question text alone.
These findings clearly indicate that our model’s ability to integrate both visual and textual modalities results in significant enhancements in retrieval accuracy. This surpasses both mere modality addition and the question-only configuration, highlighting the effectiveness of our modality fusion technique for composed retrieval tasks. Our approach provides advantages that cannot be achieved through separate handling or limited integration of text and image data.
@@ 6 Conclusion
In conclusion, our work represents a significant advancement in different downstream tasks through MLLM4PUE, a novel framework leveraging MLLMs to achieve universal multimodal embeddings. By effectively integrating image and text data, our model addresses the limitations of previous CLIP-based methods. We also introduce PMEB, which provides a standardized and comprehensive benchmark for evaluating multimodal capabilities, promoting reproducibility and comparability in the field. This study also paves the way for future research to utilize MLLM-based models to support a wide range of downstream tasks and unify the research direction for foundation models in pathology.

### v2

@@ 5 Results
5.1 Zero-shot Cross-modal Retrieval
We first evaluate our methods on the retrieval task on Arch-PubMed [12] and Arch-book datasets [12], where the query is an image or text, with the target being the corresponding paired text or image. Recall@k metric is employed to assess the model’s ability to retrieve relevant pathology image-text pairs among top k results. As shown in Table 2, our models consistently outperform the baseline models, especially at higher recall values. Notably, in the image-to-text (i2t) task on Arch-book dataset [12], our MLLM4PUE trained on Openpath dataset [15] achieves a Recall@5 of 0.185, indicating an approximate 9% improvement over PLIP [15]. Moreover, our PathCap-pretrained model surpasses the performance of PathCLIP [37] on both benchmark datasets. A similar enhancement is observed in the text-to-image (t2i) retrieval task. For instance, on the Arch-PubMed dataset [12], our model achieves a Recall@10 of 0.193, marking a 10% increase in performance. These results collectively demonstrate that our framework excels at capturing multimodal embeddings more effectively than CLIP-based models with the same training data. Furthermore, it is important to note that the E5-V model [18] yields significantly lower recall scores across all tasks and datasets due to the lack of fine-tuning with pathology-specific data.
@@ 5.1 Zero-shot Cross-modal Retrieval
Table 4: Zero-shot composed retrieval performance on the Quilt-VQA [35] and Quilt-VQA-RED [35] datasets, with Recall metrics reported at various thresholds. MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [15], PathCap [37], Quilt1M [16], respectively. Bold values indicate the best performance. Same color represents models trained on the same dataset.
5.2 Zero-shot Patch Classification
We then evaluate the performance of zero-shot classification on patch-level pathology images across 13 classification tasks from our PMEB. In this downstream task, each query is an image, and the target is an extended sentence of the class name as described in Section 3.4. Table 3 illustrates that our MLLM4PUE consistently surpasses baseline models on most datasets. Notably, with the same training data, MLLM4PUE-O outperforms PLIP [15] across all datasets. Similarly, MLLM4PUE-Q achieves the best performance compared to QuiltNet [16], demonstrating the effectiveness of our pretraining strategy on the same training dataset. While CONCH [31] attains the highest score on SkinCancer [22] and Wsss4luad [14], its superior performance stems from utilizing a much larger private dataset (1.16M). In contrast, our model, MLLM4PUE-P, outperforms all baseline models across seven datasets with only 0.2 million training samples, and is only 0.02 less than CONCH [31] on the Wsss4luad dataset [14]. This underscores the advantages of leveraging vision-language alignment in pathology representation learning. These findings collectively highlight the effectiveness of our approach, especially considering that our models were trained on fewer samples, further demonstrating their data efficiency in pathology-specific zero-shot classification. The accuracy metric shows a similar trend, where MLLM4PUE-P outperforms the baselines on most datasets.
5.3 Zero-shot Composed Retrieval
[…]
5.4 Visualization of Modality Gap
To illustrate the modality gap, we employed t-SNE visualization to compare the embeddings of our MLLM4PUE-P model with those of PathCLIP [37], both trained on the Patchcap dataset [37]. As shown in Fig 4, the qualitative analysis clearly shows that the embeddings produced by MLLM4PUE-P form tighter clusters in the low-dimensional space, whereas those from PathCLIP [37] are more dispersed. This difference reflects a significant improvement in cross-modal alignment of our method.
Figure 4: Distribution of image and text embeddings on Arch-PubMed [35] and Arch-Book [35] dataset.
To further quantify the modality gap, we adopted the metric $||\Delta||_{gap}$ as proposed in Liang et al.  [26]. Our experiments reveal that, on the Arch-PubMed dataset [35], our method reduces the gap from 0.9284 to 0.2587, and on the Arch-Book dataset [35], from 1.0408 to 0.3739. These results provide strong evidence that our model can significantly narrow the gap between image and text embeddings.
5.5 Ablation Studies
@@
Robustness of PMEB. Since some datasets in our PMEB are super large, we choose a random sample of 1,000. To assess the robustness of our benchmark, we conduct an ablation study by computing the wf1 scores using 1,000, 3,000, 5,000, and all available samples for each dataset. As shown in Figure 5, the performance change between 1,000 and the full dataset remains within a small margin across all datasets, confirming that using only 1,000 samples provides a reliable approximation of the complete dataset’s performance. Given that the performance difference is typically within 0.01 to 0.04 across different sample sizes, using 1,000 samples significantly reduces computational costs while maintaining reliable performance. This makes PMEB highly efficient for large-scale experiments without sacrificing meaningful evaluation. We also report the result of accuracy metric and the results in CONCH [31] and PathCLIP [37] model in the Supplementary Materials, which shows the same trend.
Figure 5: Ablation study on the sample numbers of our benchmark on our MLLM4PUE-P model. We computed the wF1 scores using 1,000, 3,000, 5,000, and all available samples for each dataset.
Embedding Fusion Methods. This experiment is designed to validate the quality of our benchmark and highlight the effectiveness of our MLLM4PUE in modality fusion compared to CLIP-based model in composed retrieval tasks. As shown in Table 5, MLLM4PUE outperforms QuiltNet [16] in every setting. First, comparing “only question” results, our MLLM4PUE shows significant improvement compared with QuiltNet [16], while both models demonstrate substantial performance gains when incorporating visual information via “add embedding”. This consistent pattern across both models confirms that our benchmark has been properly curated, requiring true integration of visual and textual information rather than allowing models to infer answers solely from questions.
@@
Second, the results reveal that QuiltNet [16], a CLIP-based model, shows little difference between using only the question text and integrating image and text embeddings. This indicates that CLIP’s method lacks robust modality integration, as the inclusion of visual information fails to significantly enhance its effectiveness. Conversely, our model MLLM4PUE demonstrates substantial improvements by effectively fusing image and text information. The true strength of our approach becomes evident when employing a prompt-based method, leading to significantly higher recall scores. Specifically, it achieves Recall@5 values of 0.475 on Quilt-VQA and 0.675 on Quilt-VQA-RED, representing approximately an 8% and 12% improvement over the simple addition method, respectively, and an 18% and 25% improvement over using the question text alone. These findings clearly indicate that our model’s ability to integrate both visual and textual modalities results in significant enhancements in retrieval accuracy. This surpasses both mere modality addition and the question-only configuration, highlighting the effectiveness of our modality fusion technique for composed retrieval tasks. Our approach provides advantages that cannot be achieved through separate handling or limited integration of text and image data.
@@ 6 Conclusion
In this paper, we have explored a novel paradigm for applying MLMMs to generate multimodal universal embeddings for a broad range of pathology downstream tasks. By applying summarization prompts and inserting lightweight LoRA modules, our proposed framework, MLLM4PUE, can effectively integrates image and text data, and addresses the limitations of previous CLIP-based methods. We also introduce PMEB, which provides a standardized and comprehensive benchmark for evaluating multimodal capabilities, promoting reproducibility and comparability in the field. Comprehensive experiment results across 15 diverse zero-shot retrieval and classification tasks underscore the versatility and effectiveness of our method. We anticipate that our framework will provide valuable insights into the pathology vision-language foundation models and inspire further research in this area.